OpenStudy (anonymous):

The sum of first 14 terms of 1,1.4,1.8,.... is equal to the sum of first n terms of 5.6,5.8,6.0,.... , find the value of n

OpenStudy (anonymous):

1,1.4,1.8,.... This is in AP with first term 1 and common difference 0.4

OpenStudy (anonymous):

Use sum formula of AP

OpenStudy (amoodarya):

14/2(a1+a14)=n/2(5.6+5.6+(n-1)*0.2)

OpenStudy (amoodarya):

14/2(1+1+13(.4))=n/2(5.6+5.6+(n-1)*0.2)

OpenStudy (anonymous):

i did the first A.P and its coming 50.4

OpenStudy (anonymous):

S14 =14/2 {2*1 + (14-1)0.4} = 50.4

OpenStudy (anonymous):

i am stuck here

OpenStudy (anonymous):

5.6,5.8,6.0,.... This is in AP with common difference 0.2 andfirt term 5.6

OpenStudy (anonymous):

\[50.4=\frac{ n }{ 2 }[(2*5.6)+(n-1)*0.2]\]

OpenStudy (anonymous):

stuck here

OpenStudy (anonymous):

Sn=50.4 n/2 { 2*5.6 + (n-1)0.2}=50.4 n{11 - 0.2n}=100.8

OpenStudy (anonymous):

2*5.6 its 11.2

OpenStudy (anonymous):

11n -0.2n^2 =100.8 11n -2/10 n^2 =100.8 110n -2n^2 =1008

OpenStudy (anonymous):

11.2-0.2=11

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

110n -2n^2 =1008 2n^2-110n+1008=0 n^2-55n+504=0

OpenStudy (anonymous):

how did became 110n -2n^2 =1008

OpenStudy (anonymous):